EXPLAIN. WHY L² & ONLY ONE OF COMPONENTS OF. L ARE REGARDED AS CONSTANT OF MOTION.
( (ASKED IN UNIVERSITY EXAM FOR 20 MARKS))
1
Expert's answer
2020-01-07T12:46:29-0500

We know that if any object is moving, on any path, then

velocity = "\dfrac{dx}{dt}"

so if any function is given

"y=x^2"

now , taking the differentiation with respect to x,

"\dfrac{dy }{dt}=2x"

now put x= L

so "\dfrac{dy}{dt}=2L"

Hence velocity is directly proportional to L, not to the "L^2"